### v1

# Smart Data Portfolios:

Abstract
Growing concerns about fairness, privacy, robustness, and transparency have made it a central expectation of AI governance that automated decisions be explainable by institutions and intelligible to affected parties. We introduce the Smart Data Portfolio (SDP) framework, which treats data categories as productive but risk-bearing assets, formalizing input governance as an information-risk trade-off. Within this framework, we define two portfolio-level quantities, Informational Return and Governance-Adjusted Risk, whose interaction characterizes data mixtures and generates a Governance-Efficient Frontier. Regulators shape this frontier through risk caps, admissible categories, and weight bands that translate fairness, privacy, robustness, and provenance requirements into measurable constraints on data allocation while preserving model flexibility. A telecommunications illustration shows how different AI services require distinct portfolios within a common governance structure. The framework offers a familiar portfolio logic as an input-level explanation layer suited to the large-scale deployment of AI systems.
Keywords: AI governance, input governance, training data, portfolio theory, explainability, fairness, privacy
1 Introduction
Artificial intelligence systems increasingly influence access to credit, connectivity, health, education, and public services. When a credit application is denied, or when a child is exposed to harmful recommendations, affected individuals do not ask how gradients propagated through a model. They ask which data were used, in what priority, and under whose governance. While real disputes arise at the level of inputs, existing explainability methods focus on model internals (Strumbelj2014; Lundberg2017; Koh2017), a mismatch that persists even in contemporary AI governance frameworks (Cordeiro2026). This leaves institutions without a principled way to justify which data categories were admissible, how they were used, and how those choices shaped downstream outcomes.
Current regulatory regimes reflect this tension. GDPR Article 15 requires meaningful explanations of automated decisions, the NIST AI Risk Management Framework emphasizes data and input governance (NIST2023), and Article 10 of the EU Artificial Intelligence Act mandates rigorous documentation of training, validation, and testing data (EUAIAct2024). Yet none of these instruments provide procedures for selecting, weighting, or explaining the data that feed high-stakes AI systems. The result is a structural gap between legal expectations of explainability and the technical practice of assembling training corpora. Recent system-level transparency assessments show that even state-of-the-art foundation models provide only limited disclosure of input selection and governance practices (Bommasani2025). The gap widens as AI systems become more adaptive, tool-integrated, and continuously updated in deployment.
A growing body of research points to the same underlying problem. Studies on fairness show that disparate impacts often arise from imbalanced or historically patterned data (Barocas2016; Mehrabi2021). Works on data privacy demonstrate that disclosure risk grows with granularity, linkability, and weak provenance (Narayanan2008; Rocher2019). Robustness failures similarly trace back to narrow or weakly curated training distributions (Recht2019; Taori2020). Taken together, these studies show that high-stakes AI systems require a structured approach to governing the data that underpin model behavior (Sambasivan2021).
We introduce the Smart Data Portfolio (SDP) framework, which treats data categories as productive but risk-bearing information assets whose allocation must be governed, explained, and audited. SDPs provide a dual contribution. First, they transform the qualitative process of data selection into a structured governance problem by defining two measurable quantities: Informational Return and Governance-Adjusted Risk, whose interplay generates a Governance-Efficient Frontier. Second, they supply a model-agnostic explanation layer where institutions justify decisions not by revealing internal model mechanics but by disclosing the regulated portfolio of authorized data. These are documented through Data Portfolio Statements, Data Portfolio Cards, and Consumer Portfolio Reports that record validated portfolio metrics.
This portfolio-based explanation protocol aligns with long-standing practices in financial supervision, where institutions optimize within risk envelopes defined by eligibility rules, concentration limits, and capital buffers. SDPs adopt the same supervisory logic for training data. Regulators set admissible categories, weight bands, and policy risk caps, while institutions select portfolios that maximize informational value subject to these constraints. Crucially, this supervisory structure fits naturally within existing data-governance, risk-management, and compliance processes.
Thus we make four contributions. (1) We develop Smart Data Portfolios as an explanation layer for AI systems, enabling decisions to be justified through governed input portfolios rather than model internals. (2) We formalise informational return, governance-adjusted risk, the governance-efficient frontier, and the policy risk cap, creating a quantitative foundation for input governance. (3) We propose a governance architecture with regulatory tools and reporting artefacts that support transparent and auditable deployment. (4) We illustrate the framework in the telecommunications sector, where heterogeneous applications require distinct governed portfolios under a shared regulatory structure.
The remainder of the paper is organized as follows. Section 2 reviews related research, while Section 3 defines the SDP framework and its principles. Section 4 develops the quantitative framework and derives the governance-efficient frontier. Section 5 introduces the reporting stack that operationalizes SDPs as an explanation layer. Section 6 presents the sectoral illustration, and Section 7 concludes with implications for AI governance.
2 Background and Related Literature
Research on AI governance spans a wide range of technical, legal, and institutional concerns that have largely evolved in separate strands. Technical literatures have examined fairness, privacy, robustness, and micro-level explainability largely as distinct problems, while parallel work has studied informational value and efficiency. Policy frameworks have only more recently attempted to integrate these perspectives into coherent governance principles. What remains insufficiently developed is a unifying governance approach that can deliver responsible and explainable AI in a form that is consistent, scalable, and intelligible across applications and sectors. This section reviews five strands of research that motivate our focus on input data as a governed portfolio of information assets: (1) transparency and explainability, (2) fairness and bias, (3) privacy and data protection, (4) robustness and data quality, and (5) the value of information and efficiency.
2.1 Transparency and explainability
An early response to the opacity of machine learning systems focused on explainability at the level of individual predictions. Strumbelj2014 decompose model outputs into feature-level contributions. Lundberg2017 provide a unified framework for explaining outputs based on additive feature attributions, while Koh2017 develop influence functions that trace model behavior back to particular portions of the training set.
Technical approaches to documentation and reporting have complemented these explainability efforts. Mitchell2019 introduce Model Cards as structured summaries of model behavior, including information about training data, evaluation conditions, and subgroup performance. In parallel, dataset documentation frameworks such as Datasheets for Datasets (Gebru2021) record provenance, composition, collection protocols, and intended uses; Paullada2021 survey this literature and highlight its role in improving transparency at the level of individual datasets. Taken together, these contributions provide developers, users, and regulators with clearer insight into how models behave and how particular datasets or examples influence outputs. They also emphasize the value of well-documented datasets, while leaving open the question of which data sources were permissible and under which constraints.
2.2 Fairness and bias
A large literature examines how machine learning systems can produce unequal or discriminatory outcomes. Barocas2016 analyze how data collection, feature construction, and labeling choices can embed historical inequities into ostensibly neutral models, leading to disparate impacts even in the absence of explicit use of protected attributes. Hardt2016 formalize a notion of equality of opportunity in supervised learning and show that standard training objectives can generate systematic differences in error rates across groups, which motivates explicit fairness constraints and post-processing adjustments. Mehrabi2021 provide a comprehensive survey of bias and fairness in machine learning, cataloging sources of harm that arise from coverage gaps, skewed labels, feedback effects, and mis-specified objectives.
Sectoral studies reinforce these concerns as well. In health care, for example, Chen2021 review ethical machine learning practices and document how imbalanced or unrepresentative training data can translate into clinically significant disparities in performance across patient groups.
These contributions converge on the view that fairness outcomes are tightly linked to the composition and curation of training data, not only to the choice of prediction algorithms. They also suggest that fairness metrics and constraints are most meaningfully evaluated at the level of data mixtures and subgroup coverage than solely at the level of model outputs.
2.3 Privacy and data protection
Privacy and data protection have long been central concerns in data-intensive systems. Early work on microdata anonymization showed that combinations of quasi-identifiers can sharply increase the risk of reidentification even when direct identifiers are removed (Samarati2001). Subsequent analyses of large-scale data practices demonstrate that linkage across heterogeneous sources creates cumulative disclosure risk, often in ways that are difficult to anticipate (Narayanan2008; Rocher2019). These findings underscore that privacy loss is primarily a function of data structure, granularity, and cross-system interoperability rather than of model design alone.
A complementary line of work develops formal guarantees for limiting individual information leakage. Differential privacy provides mechanisms that bound the inference risk associated with queries or model training and quantifies how privacy loss compounds under composition (Dwork2014; Abowd2018; Wood2020). This literature makes explicit that privacy exposure depends on the magnitude, sensitivity, and reuse of the underlying data, and that even well-designed models can violate privacy when trained on highly identifying inputs.
Together, these strands emphasize that privacy risk is not simply an output-level property of models. It arises from decisions about which data to collect, how long to retain them, how to link them across systems, and how to share or release them. They also imply that privacy requirements will often need to be expressed as constraints on the use and weighting of particular data categories, especially those that are highly identifying or sensitive.
Together, these strands emphasize that privacy risk is not simply an output-level property of models. It arises from decisions about which data to collect, how long to retain them, how to link them across systems, and how to share or release them. They further underscore that governing inferences and data combinations is central to data protection in contemporary AI systems (Wachter2019). As a result, privacy requirements will often need to be expressed as constraints on the use and weighting of particular data categories, especially those that are highly identifying or sensitive.
2.4 Robustness and data quality
Robustness to changes in data distributions has been the focus of another active line of work. Recht2019 construct a test set for ImageNet and show that image classifiers that perform well on the original benchmark can suffer notable accuracy declines when evaluated on their new but closely related dataset. Taori2020 extend this analysis to a broader range of natural distribution shifts and document systematic performance degradation across models and training regimes. These results suggest that models can be fragile to relatively modest changes in the data generating process, especially when they have been optimized on a narrow distribution of training examples.
In parallel, research on data quality has mapped out common sources of degradation in machine learning procedures. Jakubik2024 review the state of data quality for machine learning and identify issues such as label noise, incomplete or biased coverage, missing documentation, and the inclusion of manipulated or low-veracity sources. They argue that many observed failures of deployed systems can be traced back to shortcomings in data collection and curation rather than to model architecture. This line of work reinforces the idea that robustness and reliability are properties of the combined data and model system, particularly in settings where deployment conditions evolve over time.
2.5 Value of information and efficiency
A further strand of research considers the value and efficient use of information. Early work in decision analysis, notably Howard1966, develops information value theory and shows that the expected benefit of additional information can be quantified in terms of improved decision quality. Hubbard2014 translates these ideas into practical frameworks for measuring the value of intangibles in organizational settings, emphasizing that many institutions collect more information than they can effectively use.
In finance, Markowitz1952 introduce the idea that portfolio choice can be framed as a balance between expected return and risk, establishing diversification and risk constraints as central features of decision under uncertainty. Subsequent research develops alternative risk measures that capture tail exposures more effectively, most notably Conditional Value-at-Risk (Rockafellar2000). Related work on coherent risk measures (Artzner1999) provides axiomatic foundations for evaluating optimal portfolio allocation. Elton2014 survey the broader development of modern portfolio theory where regulated constraints and model-based estimates jointly determine feasible allocation sets.
In the context of data, Ghorbani2019 assign values to individual data points or subsets based on their marginal contribution to model performance, building on the Shapley value concept from cooperative game theory (Shapley1953). Subsequent work extend these ideas to federated learning, data marketplaces, and scalable privacy-aware settings under heterogeneous data (Wang2020; Sun2023; Tian2022; Wang2023). This line of research shows that not all data are equally valuable, and formalizes information as an asset with quantifiable marginal contributions.
2.6 Governance frameworks and emerging principles
Policy frameworks have begun to synthesize these research strands into articulated governance aims. The OECD Principles on Artificial Intelligence (OECD2019), the NIST AI Risk Management Framework (NIST2023), the EU Artificial Intelligence Act (EUAIAct2024), and the U.S. Blueprint for an AI Bill of Rights (OSTP2022) all emphasize the need for transparency and documentation of AI systems, fairness and non-discrimination in outcomes, privacy and data protection, robustness and security, and proportional use of data and computational resources. They require, to varying degrees, that high-risk AI systems document training data and model behavior, assess subgroup performance, safeguard privacy, and demonstrate robustness under realistic operating conditions.
Despite these advances, existing governance frameworks largely stop at articulating normative objectives, rather than providing a general architecture capable of sustaining them. They state what AI systems should achieve, but leave unresolved how institutions are to organize, constrain, and explain the data choices that underpin automated decisions. As a result, training data are typically assembled through inherited workflows and ad hoc judgments about value and risk, with limited capacity for consistent external scrutiny across deployments. This need for explanation and justification through input governance motivates the Smart Data Portfolio framework developed in this study.
3 The Smart Data Portfolio Framework
Building on the preceding literature, the Smart Data Portfolio framework formalizes input governance as a structured allocation problem over regulated data categories. We define a data category as a standardized class of inputs whose internal composition may change over time, but whose use is governed, constrained, and reported as a single entity. Categories are specified by regulators at an intermediate level of detail, capturing shared provenance, governance-risk characteristics, and semantic scope, rather than raw records or model-specific features. This level of abstraction enables data inputs to be governed in a way that is stable, comparable, and auditable across sectors and institutions.
An SDP represents a constrained allocation of such data categories for model training. Formally, it specifies a vector of portfolio weights that determine how much each category contributes to the overall training mixture. Changing these weights alters two fundamental quantities: informational return and governance-adjusted risk.
Informational return measures the empirically validated performance a model can extract from a given mixture of inputs. The operationalization of informational return is domain-dependent and defined through task-appropriate performance metrics and regulator-approved validation protocols. Governance-adjusted risk, on the other hand, captures the regulatory and institutional exposure generated by a given data portfolio. Unlike model performance, which reflects predictive accuracy, governance risk captures the expected burden associated with deploying a system in real-world, regulated environments. These burdens arise from systematic disparities across affected groups, weaknesses in data provenance and documentation, and instability under foreseeable changes in operating conditions.
Importantly, governance-adjusted risk is a continuous quantity rather than a binary concept. Even strong data-governance hygiene cannot eliminate governance risk. Documentation, lineage tracking, consent management, and minimization requirements ensure procedural compliance but do not equalize the fairness, privacy, provenance, or robustness properties of different data categories. That is why informational return and governance-adjusted risk cannot be collapsed into a single objective or treated as interchangeable. They must be evaluated jointly at the SDP level.
SDPs with different allocations of data categories yield different combinations of informational value and governance-adjusted risk. As a result, some portfolios are strictly superior, in that no alternative provides higher informational return at the same level of governance-adjusted risk. The set of such portfolios for each level of risk defines the Governance-Efficient Frontier. Institutions typically select portfolios on this frontier to maximize informational return subject to the governance constraints, while regulators control the aggregate risk level.
Figure 1 illustrates SDP optimization under a policy-defined risk cap. Each point on the plane corresponds to an SDP, that is, a feasible mixture of data categories. The solid curve represents the Governance-Efficient Frontier, tracing the upper envelope of attainable informational return for each level of governance-adjusted risk. Hollow markers below indicate inferior SDPs that achieve lower informational return than frontier portfolios at comparable risk. The vertical line represents the Policy Risk Cap determined by regulation. Points A and B lie within the admissible region to the left of the Risk Cap, while points C and D lie to the right of the cap and are therefore infeasible. The governance-optimal portfolio is the frontier point that maximizes informational return while remaining within the allowed risk level.
Governance-Adjusted RiskInformational Return00.20.20.40.40.60.60.80.81100.20.20.40.40.60.60.80.811Governance-Efficient FrontierPolicy Risk CapABCDGovernance-optimal portfolioinferior SDPsinfeasible
Figure 1: Illustrative governance-efficient frontier under a policy-defined risk cap. Filled markers denote admissible SDPs. Hollow markers denote feasible but dominated SDPs that lie below the frontier. Points to the right of the Risk Cap are policy-infeasible. The governance-optimal portfolio is the highest-return point on the frontier within the admissible region.
In addition to the policy risk cap discussed above, regulators can shape the design space through Admissible Data Categories as well as Governance Weight Bands. Admissible categories determine which classes of data may be used at all, while weight bands impose upper or lower bounds on their contribution to training mixtures. Together with the policy risk cap, these instruments constrain input composition without prescribing model architectures, training procedures, or optimization methods.
The portfolio-based perspective offers several conceptual advantages. First, it creates measurable governance objects where existing approaches rely largely on qualitative judgment: informational return and governance-adjusted risk become explicit quantities that can be compared across systems and over time. Second, it separates technical performance from governance burden, giving institutions and regulators a shared vocabulary for assessing how data choices shape system behavior. Third, because the logic operates at the level of data categories rather than model architectures, it provides a model-agnostic and sector-neutral input layer for governance that applies across different domains. Fourth, explainability becomes an input-level property, enabling institutions to justify and defend outcomes by referencing approved, risk-bounded data portfolios rather than opaque internal mechanics. Finally, by linking regulatory tools such as risk caps and weight bands directly to portfolio structure, SDPs allow firms and regulators to reason transparently about feasible alternatives when certain data categories must be constrained or excluded.
These ideas draw on long-established practices in financial supervision. Rather than prescribing valuation models, regulators define admissible regions through eligibility rules, concentration limits, capital buffers, and exposure constraints, while institutions optimize within these boundaries using their own internal methods. The SDP framework adopts this supervisory logic for AI inputs, formalizing how governance parameters shape the feasible design space without constraining model choice or innovation.
4 Methodology
This section formalizes the SDP framework by defining portfolio-level performance and governance risk measures, the resulting governance-efficient frontier, and the regulatory constraints that shape admissible data allocations.
4.1 Informational Return
Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ denote the available data categories, and let $w=(w_{1},\ldots,w_{n})$ be a nonnegative weight vector with $\sum_{i}w_{i}=1$. Training a model from class $\mathcal{M}$ on the mixture induced by $w$ yields a task-specific performance metric $M(w;\mathcal{M})$, evaluated on an institutionally approved validation procedure. Depending on the application, this may correspond to AUC, cross-entropy loss, RMSE, calibration error, or other sector-appropriate performance indicators.
We define the informational return of the portfolio as
$\mu_{p}(w)=M(w;\mathcal{M}),$ (1)
which captures the empirically validated performance that the institution’s modeling pipeline extracts from the selected mixture of data categories. Informational return is intentionally defined relative to the modeling class $\mathcal{M}$ as it is not an intrinsic property of the data, but reflects how effectively the institution’s chosen models and training procedures convert inputs into predictive performance.
In practice, institutions estimate $\mu_{p}(w)$ using standard validation protocols already embedded in their ML pipelines, such as cross-validation, temporal splits, or regulator-approved holdout procedures. The SDP framework does not prescribe a particular metric or estimation method, requiring only that informational return be computed consistently and documented in a form suitable for regulatory review.
This treatment mirrors modern portfolio theory, where the efficient frontier is defined relative to an institution’s return model and estimation choices rather than as an absolute property of the assets themselves (Markowitz1952). Accordingly, different institutions may trace different governance-efficient frontiers even under identical regulatory constraints, reflecting heterogeneous modeling strategies rather than governance failures.
4.2 Governance-Adjusted Risk
Governance-adjusted risk is a composite measure that captures expected governance burdens. These burdens typically arise from fairness dispersion, documentation and provenance defects, or operational instability. Each of these dimensions is explicitly recognized in contemporary AI governance regimes, including the EU Artificial Intelligence Act (EUAIAct2024), the OECD AI Principles (OECD2019), and the NIST AI Risk Management Framework (NIST2023). To operationalize these concerns at the input level, we decompose governance-adjusted risk into three components:
  1. 1.
    Fairness Dispersion ($F(w;\mathcal{M})$). Variation in model performance across relevant subgroups, measured through gaps in error rates, acceptance rates, or domain specific fairness indicators (Barocas2016; Hardt2016). As with informational return, this is evaluated on the trained model.
  2. 2.
    Provenance and Quality Defect Score ($P(w)$). Documentation and lineage risks such as missing consent metadata, uncertain licensing, or high label noise (Gebru2021). This component is model agnostic, reflecting properties of the underlying data lineage rather than model behavior.
  3. 3.
    Robustness Volatility ($R(w;\mathcal{M})$). Instability under benign distribution shifts, estimated from out-of-time or out-of-subset tests (Recht2019; Taori2020). One operational definition is
    $R(w;\mathcal{M})=\max_{s\in\mathcal{S}}\left|M_{\mathrm{val}}(w;\mathcal{M})-M_{\mathrm{shift}\;s}(w;\mathcal{M})\right|,$ (2)
    where $\mathcal{S}$ is a regulator-specified set of benign shifts such as temporal drift, mild covariate corruption, or geographic reweighting. This captures operational fragility that can generate governance harm.
For comparability across institutions and to eliminate scale ambiguity, each component is mapped to a common unit interval using regulator-defined transformations. As a result, $F(w;\mathcal{M}),P(w),R(w;\mathcal{M})\in[0,1]$, and the relative magnitudes of the policy parameters $(\alpha,\beta,\gamma)$ reflect regulatory priorities rather than arbitrary metric scales.
While nonlinear interactions are possible, governance-adjusted risk is defined linearly for regulatory interpretability as follows:
$\sigma_{p}(w)=\alpha F(w;\mathcal{M})+\beta P(w)+\gamma R(w;\mathcal{M}),$ (3)
In practice, regulators calibrate the parameters $(\alpha,\beta,\gamma)$ by mapping each component to its expected institutional cost under applicable law and sectoral penalties. The schedule need not be universal; only stability within a regulatory cycle is required. Continuity of each component implies that the attainable region in the information-risk plane is compact, ensuring that the efficient frontier exists.
The composite quantity $\sigma_{p}(w)$ treats governance burdens as an average-case measure. In high-stakes settings, however, regulators are primarily concerned with tail exposure, which refers to rare but severe fairness, privacy, or provenance failures. Consequently, a natural extension replaces $\sigma_{p}(w)$ with its Conditional Value-at-Risk (CVaR) analogue (Rockafellar2000).
Let $L(w)=\alpha F(w;\mathcal{M})+\beta P(w)+\gamma R(w;\mathcal{M})$ denote the composite governance loss and let $\operatorname{ES}_{\eta}$ denote Expected Shortfall at tail probability $\eta\in(0,1)$. The tail-sensitive governance risk is defined as:
$\sigma^{\text{CVaR}}_{p}(w)=\operatorname{ES}_{\eta}\!\bigl[L(w)\bigr]=\frac{1}{\eta}\int_{0}^{\eta}\operatorname{VaR}_{u}\!\bigl[L(w)\bigr]\,du.$ (4)
Following Rockafellar2000, CVaR admits a convex reformulation via auxiliary variables, enabling implementation through linear programming. Regulators may therefore replace the policy risk cap with $\sigma^{\text{CVaR}}_{p}(w)\leq\bar{\sigma}$ when tail control is required. This refinement provides a more advanced risk measure aligned with contemporary supervisory practice in financial regulation.
4.3 Governance-Efficient Frontier
Each portfolio $w$ generates a point $\bigl(\mu_{p}(w),\sigma_{p}(w)\bigr)$ in the information-risk plane. Varying the weight vector over the simplex defines the attainable region. The governance-efficient frontier is the upper envelope of attainable informational return at each level of governance risk.
In practice, institutions construct the frontier empirically. Starting from a regulator-approved feasible region $\mathcal{W}\subseteq\Delta^{n-1}$, the procedure can be summarized as follows:
  1. 1.
    Sample candidate portfolios $w^{(k)}\in\mathcal{W}$ using grid search, Dirichlet sampling, or Bayesian exploration.
  2. 2.
    For each candidate, train or fine tune the model class $\mathcal{M}$ on the induced mixture and evaluate the validation metric and governance components $F(w^{(k)};\mathcal{M})$, $P(w^{(k)})$, and $R(w^{(k)};\mathcal{M})$.
  3. 3.
    Compute the portfolio-level quantities
    $\mu_{p}^{(k)}=M(w^{(k)};\mathcal{M}),\qquad\sigma_{p}^{(k)}=\alpha F(w^{(k)};\mathcal{M})+\beta P(w^{(k)})+\gamma R(w^{(k)};\mathcal{M}).$ (5)
  4. 4.
    Discard dominated points and fit the upper envelope of $\{(\sigma_{p}^{(k)},\mu_{p}^{(k)})\}$ by concave hull or smooth interpolation to obtain the governance-efficient frontier.
Because informational return and governance-adjusted risk are not statistically dual quantities, the shape of the frontier in $(\sigma_{p}(w),\mu_{p}(w))$ space is empirical. Existing studies, summarized in Section 2, suggests that in many regulated consumer-facing settings the frontier is approximately concave or inverse-U shaped: marginal informational gains saturate once core categories are included (Recht2019; Taori2020), while fairness, privacy, and provenance risks typically rise with additional behavioral or weakly curated inputs (Mehrabi2021; Sambasivan2021). A detailed empirical analysis of frontier shapes and associated confidence intervals is left for future work.
4.4 Portfolio optimization and regulation
Once the governance-efficient frontier is established, portfolio optimization proceeds by selecting the data allocation that maximizes validated performance subject to regulator-defined governance constraints. The central regulatory instrument is the policy risk cap, which sets an upper bound on aggregate governance exposure while preserving institutional discretion over model architectures, training procedures, and validation strategies.
For a given modeling class $\mathcal{M}$ and policy risk cap $\bar{\sigma}$, the optimization problem can be written as
$\max_{w}\;M(w;\mathcal{M})\qquad\text{s.t.}\qquad\sigma_{p}(w)\leq\bar{\sigma},\;\sum_{i}w_{i}=1,\;w_{i}\geq 0.$ (6)
This formulation mirrors supervisory practice in financial regulation, where institutions optimize within a regulator-defined risk envelope rather than complying with prescriptive design rules.
In practice, optimization is carried out over a constrained feasible region defined jointly by the policy risk cap and additional category-level governance instruments. Because neighboring portfolios correspond to nearby weight vectors, warm starts and incremental fine-tuning substantially reduce computational cost. For high-dimensional or expensive training regimes, Bayesian optimization with Gaussian Process surrogates provides an efficient exploration strategy (Snoek2012; Frazier2018).
Admissible data categories
In many applications, sectoral or jurisdictional governance frameworks impose explicit eligibility rules on the types of data that may be used in AI systems. Within the SDP framework, these rules take the form of Admissible Data Categories, which define which classes of data are permitted to enter a portfolio at all. Eligibility reflects legal, ethical, and sector-specific norms related to privacy, documentation, consent, and purpose limitation.
Formally, admissibility constraints operate by excluding specific categories or aggregated groups from the portfolio weight vector $w$. They define the outer boundary of the feasible region prior to any frontier-based optimization, ensuring that subsequent portfolio selection occurs only within regulator-approved informational boundaries. Typical examples include outright prohibitions or hard limits on high-risk data classes:
$\sum_{i\in H}w_{i}=0\qquad\text{or}\qquad\sum_{i\in H}w_{i}\leq\beta.$ (7)
Such eligibility constraints operate independently of model performance. Even if a prohibited category would yield high informational return, it remains excluded from consideration.
Governance weight bands
While admissible data categories specify which data may be used, Governance Weight Bands determine how much of each admissible category may contribute to a portfolio. These further translate qualitative governance objectives into quantitative lower or upper bounds on category-level allocations, playing a role analogous to concentration limits or exposure caps in financial regulation.
Let $L_{i}$ and $U_{i}$ denote regulator-defined bounds for category $i$. A governance weight band takes the form
$L_{i}\leq w_{i}\leq U_{i},$ (8)
with $0\leq L_{i}\leq U_{i}\leq 1$. Lower bounds can ensure baseline representation of low-risk or high-integrity data sources, while upper bounds can limit reliance on sensitive, weakly documented, or behaviorally intrusive categories.
Because weight bands act directly on the allocation vector $w$, they shape the feasible region prior to optimization and prevent circumvention through internal regrouping of data sources. They can also be imposed on aggregated category groups. For example, partitioning categories into low-risk registries $\mathcal{R}$, operational metrics $\mathcal{O}$, and behavioral traces $\mathcal{B}$, a regulator may specify:
$\sum_{i\in\mathcal{R}}w_{i}\geq L_{\mathcal{R}},\qquad\sum_{i\in\mathcal{O}}w_{i}\in[L_{\mathcal{O}},U_{\mathcal{O}}],\qquad\sum_{i\in\mathcal{B}}w_{i}\leq U_{\mathcal{B}}.$ (9)
These constraints ensure minimum reliance on stable data sources, balanced use of operational indicators, and capped exposure to high-risk behavioral data.
Interaction with the policy risk cap
Eligibility rules and governance weight bands operate jointly with the policy risk cap. While admissibility and weight bands constrain where portfolios may lie in allocation space, the policy risk cap constrains how much aggregate governance burden a portfolio may carry. Together, these instruments define a regulated feasible region within which institutions select the governance-optimal portfolio on the efficient frontier.
This separation of roles allows regulators to enforce high-level governance objectives while leaving institutions free to optimize performance within clearly articulated and auditable boundaries.
4.5 Attribution and interpretability
Although the frontier uses the aggregate quantities $(\mu_{p}(w),\sigma_{p}(w))$, institutions may require transparency regarding the contribution of individual data categories. Because $M(w)$ is generally nonlinear in $w$, additive decompositions of the form $\mu_{p}\approx\sum_{i}w_{i}\mu_{i}$ lack theoretical justification. As a result, a cooperative game theoretic formulation is required.
Let $\mathcal{N}=\{1,\ldots,n\}$ be the set of categories and let $M(S)$ denote performance when training only on categories $S\subseteq\mathcal{N}$. The Shapley value for category $i$ (Shapley1953; Strumbelj2014; Ghorbani2019) is given as:
$\phi_{i}=\sum_{S\subseteq\mathcal{N}\setminus\{i\}}\frac{|S|!\,(n-|S|-1)!}{n!}\big(M(S\cup\{i\})-M(S)\big).$ (10)
Shapley values provide transparent attribution without relying on model specific heuristics. Importantly, they do not influence frontier construction, which depends solely on aggregate portfolio level quantities. They instead serve a justificatory role, supporting explanation and contestation of portfolio choices without altering governance constraints or optimization outcomes.
While exact computation is exponential in $n$, practical estimators for $\phi_{i}$ based on Monte Carlo sampling, influence functions, or gradient based methods are well established (Koh2017). For portfolios with a moderate number of categories, sampled permutations on the order of a few hundred to a few thousand often yield stable estimates in empirical studies (Jia2019).
5 Policy and Implementation
The Smart Data Portfolio (SDP) framework is designed as an input-level governance architecture rather than a model-level control mechanism. Its central policy contribution is to make data-selection decisions explicit, measurable, and auditable, while preserving institutional discretion over model choice, validation procedures, and optimization routines. By operating at the level of data allocation, SDPs translate high-level governance objectives into enforceable constraints without prescribing technical implementations, consistent with emerging standards such as ISO/IEC 42001 (ISO42001). This practical focus is especially important for agentic AI systems, whose adaptive and multi-step decision processes can render output-level explanations unstable over time.
Institutional scope and primary users
The framework is most immediately actionable for three groups. First, sector regulators and supervisory agencies face the task of translating broad statutory mandates into auditable, operational rules for high-risk AI systems. Second, specialised AI-governance consultancies increasingly require quantitative, defensible artifacts to support compliance, auditing, and advisory services across jurisdictions. Third, in-house risk, compliance, and data-governance teams in data-intensive firms bear direct accountability for data-selection decisions and must justify those choices to regulators, auditors, and senior management.
These actors share three structural characteristics: direct responsibility for data governance outcomes, authority to impose portfolio constraints, and a practical need for model-agnostic tooling. By framing data allocation as a portfolio-optimization problem, SDPs provide these stakeholders with a common vocabulary and a credible pathway for implementing transparency, fairness, privacy, and robustness requirements across heterogeneous AI systems.
Supervisory workflow and portfolio iteration
Policy implementation under the SDP framework proceeds iteratively. Regulators specify governance parameters, including admissible data categories, governance weight bands, and an overall policy risk cap. Institutions then construct candidate data portfolios and evaluate them using the informational and governance-risk metrics defined in Section 4. Within these constraints, institutions retain full discretion over model class, training procedures, and validation strategies, but must justify resulting input allocations in a structured and reproducible manner.
Supervisory oversight occurs through periodic review of submitted portfolios. When constraints are violated or governance thresholds exceeded, regulators may require rebalancing or impose tighter bounds. As data environments evolve through concept drift, the introduction of new data sources, or changes in legal requirements, institutions revisit the portfolio-optimization problem and adjust weights accordingly. This cyclical structure mirrors established practices in financial supervision and model-risk governance, adapted to the specific challenges posed by data-centric AI systems.
Reporting artifacts and governance objects
To support this workflow, we propose three complementary reporting artifacts that together form an auditable disclosure trail. These artifacts can be generated using existing workflows, without requiring new technical infrastructure or model instrumentation.
The Data Portfolio Statement (DPS) provides a public-facing summary of admissible data categories, applicable governance weight bands, and the principles governing input allocation. The Data Portfolio Card (DPC) serves as a regulatory filing that records realised category-level weights together with validated values of informational return $\mu_{p}(w)$ and governance-adjusted risk $\sigma_{p}(w)$, supported by documentation of data sources, lineage, and validation procedures. The Consumer Portfolio Report (CPR) provides consumer-facing explanations of adverse actions or decisions by referencing the approved input portfolio rather than internal model mechanics.
Taken together, these artifacts parallel disclosure regimes used in credit reporting, financial stress testing, and data-protection impact assessments, while being specifically tailored to the governance of AI training data.
Relation to existing documentation tools
Existing documentation tools, most notably Model Cards and Datasheets (Mitchell2019; Gebru2021), improve transparency by describing model behavior and individual dataset properties. However, they do not govern how multiple data sources are combined, weighted, or constrained within a training mixture. Smart Data Portfolios operate at this missing allocation layer. While Model Cards summarize model performance and Datasheets document dataset lineage and quality, SDPs provide portfolio-level explanations through standardized reporting artifacts. As AI systems increasingly mediate regulated services, this input-governance layer becomes the primary layer of accountability and explainability.
Table 1 summarises the core policy instruments within the SDP framework. Each instrument functions at the portfolio level and is designed to complement, rather than replace, existing documentation practices.
Table 1: Policy instruments of the Smart Data Portfolio framework.
Instrument Function within the framework
Admissible data categories Defines which data classes may enter portfolios, reflecting privacy, documentation, and sectoral norms.
Governance weight bands Sets quantitative minimum or maximum allocations for selected categories, analogous to concentration limits or exposure caps in financial regulation.
Policy Risk Cap Establishes an upper bound on governance burden, creating a feasible region for optimization.
Data Portfolio Statement Public summary of category-level weight ranges and data sources.
Data Portfolio Card Filed with regulators; includes validated $\mu_{p}$, $\sigma_{p}$, and detailed lineage documentation.
Consumer Portfolio Report Provides consumer-facing explanations of decisions based on approved portfolios.
Beyond these core procedures, the SDP framework can further incorporate a number of supervisory mechanisms informed by long-standing financial-risk governance. These include concentration limits for interconnected data sources, scenario-based stress testing, as well as risk-weighted exposure summaries. Such SDP extensions do not alter the structure of the underlying optimization problem. Appendix A outlines this broader regulatory design space.
6 Sectoral Illustration: Telecom as a Multi-Service AI Platform
Telecommunications operators provide an excellent test case for Smart Data Portfolios because they operate multi-service AI platforms built on heterogeneous, high-volume, and governance-sensitive data ecosystems. Privacy obligations arise under general data-protection regimes such as the GDPR, including data-minimization and purpose-limitation principles, as well as under sector-specific telecom regulations, including customer proprietary network information (CPNI) rules in the United States and parallel constraints under European data-protection and ePrivacy frameworks (FCC2023; EDPB2018). These regulatory obligations align closely with governance concerns emphasized in the academic literature (Acquisti2016; Barocas2016; Wachter2019; Recht2019; Mehrabi2021; Jakubik2024). Yet in telecom practice, these requirements remain largely procedural, offering limited guidance on how heterogeneous data sources should be selected, weighted, and constrained within AI training pipelines.
6.1 Data categories and governance profiles
Telecom operators maintain a rich set of data categories that serve different AI applications, including: (i) network KPIs such as latency, jitter, and throughput; (ii) tariff and billing records; (iii) device characteristics; (iv) coarse or granular usage aggregates; (v) location and mobility traces; and (vi) survey-based quality-of-experience (QoE) indicators. These categories differ in informational value and governance risk. Network KPIs and tariff records are structured, high-quality, and low risk. Behavioral telemetry and fine-grained location sequences are informationally rich, but also privacy sensitive, high variance, and susceptible to documentation or provenance defects.
A portfolio treatment clarifies these distinctions by mapping each category to contributions in informational return and governance-adjusted risk. Because telecom operators support AI across multiple product lines, a single governance envelope is insufficient. Each AI deployment has a distinct fairness, privacy, documentation, and robustness profile. The Smart Data Portfolio framework assigns each use case its own admissible region and governance-efficient portfolio.
6.2 Representative AI use cases
We illustrate this heterogeneity through three representative AI use cases with distinct governance requirements: device finance scoring, personalization, and network QoS prediction. These stylized examples demonstrate the mechanics of portfolio construction and the governance logic with no empirical claims about real-world performance.
High-risk use case: Device finance and credit scoring
Telecom operators routinely offer handset installment plans and small credit products, placing these models at the intersection of algorithmic fairness requirements and financial regulation (Barocas2016; Hardt2016). The privacy literature also highlights risks associated with linking behavioral data to financial decisions (Acquisti2016; Wachter2019). Regulators therefore impose tight governance constraints. In portfolio form, which can be expressed as:
$F(w;\mathcal{M})\leq F_{\max},\qquad P(w)\leq P_{\max},\qquad w_{\text{behavioral}}\leq 0.10.$ (11)
High-value, low-risk categories typically include payment and billing history, tariff selections, and device attributes. These provide strong predictive signals with limited fairness or privacy exposure. High-risk signals such as granular app telemetry or precise location sequences receive minimal allocation due to fairness and privacy exposure. The governance-efficient frontier identifies the highest attainable informational return consistent with these constraints.
This can be illustrated with a stylized example containing a single high-risk category within an otherwise balanced portfolio. The category-level values shown below represent component scores $F_{i},P_{i},R_{i}$, and their unweighted sum is reported as a Raw risk score; portfolio-level quantities are computed as weighted sums under $w$:
$F(w)=\sum_{i}w_{i}F_{i},\qquad P(w)=\sum_{i}w_{i}P_{i},\qquad R(w)=\sum_{i}w_{i}R_{i}.$ (12)
We consider five data categories ($i=1,\ldots,5$) normalized to unit weights. Table 2 shows these along with their component risk score $\sigma_{i}$.
Table 2: Data categories for portfolio example with a high-risk category
Category $w_{i}$ $F_{i}$ $P_{i}$ $R_{i}$ Raw risk score Note
Payment/billing history 0.40 0.02 0.01 0.03 0.06 Low risk
Tariff selections 0.25 0.01 0.01 0.02 0.04 Low risk
Device attributes 0.15 0.02 0.01 0.02 0.05 Low risk
Coarse usage patterns 0.15 0.04 0.03 0.05 0.12 Medium risk
Behavioral traces 0.05 0.07 0.06 0.08 0.21 High risk
Assume policy parameters $\alpha=\beta=\gamma=1/3$. Then the governance-adjusted risk of the portfolio is:
$\sigma_{p}(w)=\tfrac{1}{3}F(w)+\tfrac{1}{3}P(w)+\tfrac{1}{3}R(w).$
Aggregating across categories yields an overall risk of approximately $\sigma_{p}(w)=0.081$, which falls below a hypothetical policy risk cap of $\bar{\sigma}=0.10$. At the same time, the behavioral category weight constraint $w_{\text{behavioral}}\leq 0.10$ binds. A neighboring candidate portfolio with $w_{\text{behavioral}}=0.12$ would achieve slightly higher informational return but would violate governance constraints and is therefore ineligible. This demonstrates how governance constraints reshape the attainable frontier by excluding portfolios that rely too heavily on sensitive or weakly curated signals.
The Data Portfolio Statement (DPS) summarizes weight bands and public category-level rationales, while the Consumer Portfolio Report (CPR) provides adverse-action explanations consistent with explainability principles.
Medium-risk use case: Personalization and recommendation systems
Many telecom operators support recommendation services through video, music, social-media bundles, and personalized offers. These applications raise significant privacy and provenance concerns. Behavioral telemetry is often sensitive and disproportionately revealing (Acquisti2016).
Because contestable explanations are essential in consumer-facing services, particularly where recommendations may affect vulnerable users, a regulator may impose governance weight bands such as:
$w_{\text{behavioral}}\leq 0.15,\qquad w_{\text{QoS}}\geq 0.20,\qquad P(w)\leq P_{\max}.$ (13)
Here, governance risk is dominated by privacy sensitivity and robustness volatility, given evidence of instability in content recommendation and multimodal systems under distribution shift (Recht2019; Taori2020). The governance-efficient portfolio therefore emphasizes coarse usage aggregates, network KPIs, and minimally sensitive survey or preference data.
The resulting medium-risk portfolio can be illustrated with a stylized example shown in Table 3. As above, the table entries represent component scores $F_{i},P_{i},R_{i}$.
Table 3: Data categories for portfolio example with a medium-risk category
Category $w_{i}$ $F_{i}$ $P_{i}$ $R_{i}$ Raw risk score Note
QoS KPIs 0.25 0.01 0.01 0.03 0.05 Required minimum
Survey preferences 0.20 0.02 0.01 0.02 0.05 Low risk
Coarse usage aggregates 0.25 0.04 0.03 0.05 0.10 Medium risk
Behavioral traces 0.15 0.06 0.06 0.07 0.19 Weight cap binds
Contextual metadata 0.15 0.03 0.02 0.03 0.08 Medium risk
Assuming the same policy parameters as above, the aggregate risk is $\sigma_{p}(w)\approx 0.076$, below the hypothetical $\bar{\sigma}=0.10$, and the behavioral cap binds at $w_{\text{behavioral}}=0.15$. Compared to the device-finance example, the constraints are milder, which can lead to a governance-optimal portfolio at a relatively close point to the unconstrained frontier. The Data Portfolio Card (DPC) records weight bands, tested values of $\mu_{p}(w)$ and $\sigma_{p}(w)$, and documentation of consent, provenance, and lineage quality.
Low-risk use case: Network QoS prediction and hazard models
Network operators rely on AI to predict congestion, outages, and quality degradation. These models primarily use engineering signals and operational logs, which carry lower governance risks relative to consumer-level data. Because the data are mainly operational and engineering-driven, governance concerns relate mostly to reliability rather than distributional fairness (Recht2019; Jakubik2024).
In this case, regulators may impose light constraints such as:
$w_{\text{network\_KPIs}}\geq 0.40,\qquad R(w;\mathcal{M})\leq R_{\max}.$ (14)
Because governance risk is low, the governance-efficient portfolio closely approximates the unconstrained efficient frontier. This contrast illustrates the flexibility of SDPs, which can both tighten or relax constraints depending on use-case risk without altering the underlying optimization structure.
6.3 Comparing Portfolio Implications Across Use Cases
The three cases discussed above illustrate a natural risk gradient across telecom AI deployments: financial decisions impose the tightest governance bands, personalized services require moderate constraints, and engineering applications operate under lighter oversight. Table 4 summarizes how distinct governance priorities translate into different portfolio structures.
Table 4: Governance profiles and data portfolio implications across telecom AI use cases.
Use case Governance priorities Portfolio implications
Device finance / credit scoring Fairness, privacy, auditability Prioritize payment histories, tariffs, device data; minimal behavioral telemetry; tight policy risk cap.
Personalization / recommendations Privacy, child safety, explainability Limit granular telemetry; emphasize QoS aggregates and coarse usage groups; behavioral weight cap binds.
Network QoS prediction Robustness, provenance, operational reliability High allocation to engineering logs and KPIs; fairness constraints minimal; near-unconstrained frontier optimum.
As the table shows, SDPs can provide the telecom sector with increased governance discipline analogous to the well established financial supervision frameworks. They can also improve explainability by bridging the gap between high-level regulatory principles and operational data selection. Although illustrated in telecom, the same governance logic extends to any domain where heterogeneous data categories carry differentiated risk, including finance, insurance, health, logistics, and mobility services.
7 Conclusion
Artificial intelligence systems increasingly mediate access to credit, connectivity, health, education, and public services. Yet prevailing governance approaches continue to focus on models and outputs, while the selection and justification of training data remain largely heuristic. The Smart Data Portfolio (SDP) framework addresses this structural gap by providing a model-agnostic, input-level governance and explainability layer, translating abstract regulatory requirements into measurable and auditable constraints on data allocation, using a portfolio logic familiar from financial supervision.
SDPs quantify input governance as an explicit information-risk trade-off. Each governed data category contributes jointly to informational return and to governance-adjusted risk, and varying portfolio weights traces an attainable region whose upper envelope defines the governance-efficient frontier. Regulators shape this region by specifying admissible categories, governance weight bands, and a policy risk cap, while institutions optimize within these boundaries using their preferred models and validation procedures. These choices are disclosed through Data Portfolio Statements, Data Portfolio Cards, and Consumer Portfolio Reports, which translate portfolio-level decisions into auditable explanations for regulators and intelligible explanations for affected individuals. By separating model design from input governance, SDPs enable stable, model-agnostic accountability aligned with long-standing supervisory practice in financial regulation.
The Smart Data Portfolio framework is designed for immediate use by actors who already bear responsibility for AI governance outcomes, including sector regulators, supervisory agencies, AI-governance consultancies, and in-house risk, compliance, and data-governance teams. It can be operationalized within current governance and MLOps workflows using existing data inventories, validation results, and documentation practices, without requiring new technical infrastructure or changes to model architectures. Even coarse category definitions and approximate governance scores are sufficient to identify dominated data allocations, make performance-risk trade-offs explicit, and prioritize remediation efforts. In this sense, SDPs function as a practical diagnostic and control instrument in the short term, while remaining fully compatible with future standardization and formal regulatory adoption.
The telecom illustration demonstrates how SDPs accommodate heterogeneous applications such as device finance, personalized recommendations, and network-quality prediction within a single governance architecture. Each use case is assigned its own governed data mixture while preserving a common optimization logic and reporting grammar. Across applications, SDPs allow regulators to evaluate input choices without inspecting proprietary model architectures, clarifying how governance risks accumulate and how institutions justify decisions through input-level explanations rather than internal model mechanics.
The policy implications are immediate. Article 10 of the EU AI Act on data governance already requires that training, validation, and testing data for high-risk systems satisfy rigorously documented quality and representativeness criteria (EUAIAct2024). SDP concepts can be incorporated into this regime through sectoral data taxonomies, reference risk-weight schedules, and governance weight bands defined in implementing acts, harmonized standards, or codes of practice (OECD2019; NIST2023). In the United States, the NIST AI Risk Management Framework highlights data and input management as a core dimension of trustworthy AI, and sector regulators such as the FCC, Ofcom, and BEREC are beginning to scrutinize AI-mediated communications and automated decision systems in telecom and adjacent markets (FCC2023; EDPB2018). Requiring Data Portfolio Statements and Data Portfolio Cards alongside existing model documentation would establish a common governance vocabulary that links high-level principles to enforceable input-selection rules.
Smart Data Portfolios provide a pathway to coherent public accountability by reframing explainability as an input-level property, allowing institutions and regulators to justify automated decisions through governed data allocations rather than opaque internal model mechanics. Like any portfolio model, SDPs rely on metric choice and category delineation. As a result, future work should therefore focus on pilot implementations, empirical estimation of governance-efficient frontiers, and robustness analysis under realistic institutional constraints. Overall, SDPs show that input governance can be structured as a quantitative, optimization-based discipline that scales across sectors, complements model-level tools, and supports transparent and accountable AI deployment.
Appendix A: Finance-Inspired Extensions of Smart Data Portfolios
This appendix outlines several supervisory mechanisms that can be integrated into Smart Data Portfolios without altering the underlying optimization structure. Each mechanism has a direct analogue in long-standing financial-risk governance and provides regulators with additional levers for shaping data allocation. They are intentionally illustrative rather than exhaustive and are not required for baseline SDP implementation.
A.1 Concentration limits for connected data sources
Supervisory regimes often restrict concentrated exposure to a single counterparty. An analogous constraint for training data prevents excessive reliance on a single vendor, broker, or provenance chain. For each equivalence class $G_{k}$ of connected data suppliers,
$\sum_{i\in G_{k}}w_{i}\;\leq\;\theta_{k},$ (15)
where $\theta_{k}$ is regulator-set. These linear constraints complement the weight-band structure in the main text and guard against systemic dependence on high-risk intermediaries (BaselLE2023).
A.2 Scenario-based governance stress testing
In analogy with supervisory stress testing (e.g. DFAST), regulators may evaluate portfolios under a set of hypothetical or historically informed governance scenarios $s\in\mathcal{S}$, such as demographic shifts, sensor failures, or abrupt changes in privacy law:
$(\mu^{(s)}_{p}(w),\sigma^{(s)}_{p}(w)).$ (16)
A stress-based admissibility condition requires $\sigma^{(s)}_{p}(w)\leq\bar{\sigma}^{(s)}$ for all $s\in\mathcal{S}$. This approach exposes latent fragility in ways not captured by average-case risk measures (DFAST2024).
A.3 Risk-weighted data exposures (D-RWA)
Supervisors may summarise portfolio exposure using a single linear index analogous to bank risk-weighted assets (RWA):
$\operatorname{D\!-\!RWA}(w)=\sum_{i}\rho_{i}\,w_{i},$ (17)
where $\rho_{i}$ is a regulator-defined risk weight calibrated to fairness exposure, provenance uncertainty, or privacy sensitivity. A constraint $\operatorname{D\!-\!RWA}(w)\leq B$ provides an interpretable, scalable tool for sector-wide monitoring.
A.4 Governance baselines via risk-parity weights
A natural baseline for benchmarking institutional portfolios is an equal-risk-contribution allocation (Maillard2010), defined as the solution to
$w_{i}\,\frac{\partial\sigma_{p}}{\partial w_{i}}=w_{j}\,\frac{\partial\sigma_{p}}{\partial w_{j}},\qquad\forall i,j.$ (18)
Risk-parity weights provide a regulator-friendly neutral reference against which deviations can be audited or justified.
These extensions demonstrate the compatibility of Smart Data Portfolios with established supervisory traditions while preserving the framework’s generality and transparency.